OpenStudy (anonymous):

What could you do to eliminate the fractions in the equation c-ninths minus two-ninths equals seven-ninths? Add two-ninths to both sides of the equation. Multiply both sides of the equation by one-ninth. Subtract two-ninths from both sides of the equation. Multiply both sides of the equation by 9.

OpenStudy (jdoe0001):

c-ninths?

OpenStudy (jdoe0001):

hmmm

OpenStudy (theeric):

Hi! Since it's multiple choice, you can see which one works! Otherwise, you should do the algebra! So, do you have any thoughts? @jdoe0001 , it's probably a variable, or a constant. Either way, it doesn't change the answer! Just makes it look tougher than it is, like letters sometimes do!

OpenStudy (jdoe0001):

\(\bf \color{blue}{\cfrac{c}{9}-\cfrac{2}{9}=\cfrac{7}{9}}\\ b \times \cfrac{a}{b} = \cfrac{\square?}{\square?}\)

OpenStudy (anonymous):

im confused

OpenStudy (jdoe0001):

\(\bf \textit{what's the product of say } 3 \times \cfrac{c}{3} \ \ ?\)

OpenStudy (theeric):

@jdoe0001 is using constant numbers \(a\) and \(b\) to communicate a number relationship. jdoe0001 is speaking math :)

OpenStudy (jdoe0001):

hehe

OpenStudy (theeric):

And, if you can solve that, it'll be helpful.. Hint - if you learn what jdoe0001 is saying, try to apply it to your problem :)

OpenStudy (theeric):

Do you know where to start?

OpenStudy (anonymous):

i have to multipy each side by 9

OpenStudy (theeric):

Alright! And then there are no fractions?

OpenStudy (jdoe0001):

\(\bf b \times \cfrac{a}{b} \implies \cfrac{b}{1}\times \cfrac{a}{b} \implies \cfrac{b\times a}{1\times b} \implies \cfrac{b}{b} \times \cfrac{a}{1} \implies 1\times\cfrac{a}{1} \implies a\)